radodls.org/assignment/1936186696/assessment and the arm span, rounded to the nearest centimeter, of each person in a random sample of 10 seniors at a high school. A scatterplot of arm span versus height for the 10 seniors is shown 164 T 163 2 162 161 E 160 E 159 2158 157 156 155 t + 155 156 157 158 159 160 161 162 163 164 Height, centimeters Part A: Based on the scatterplot, descnibe the relationship between arm span and height for the sample of 10 seniors (3 points) Part B: Let x represent height, in centimeters; and let y represent arm span, in centimeters. Two scatterplots of the same data are shown. Graph 1 shows the data with the least-squares regression line y 64.494 0.5984% graph 2 shows the data with the line yx Graph 1 164 163 162 E 161 160 ch End PrtScn
coloradodls.org/assignment/1936186696/assessment Graph 1 164 T 163 句162 E 161 160 t 158 157 156 155 155 156 157 158 159 160 161 162 163 164 Height, centimeters Graph 2 164エ 162 161 160 E 159 158 157 156 d) PrtScn Home
.coloradodls.org/assignment/1936186696/assessment n 158 E 157 156 155 155 156 157 158 159 160 161 162 163 164 Height, centimeters The criteria described in the table can be used to classify people into one of three body shape categories: square, tall rectangle, or short rectangle. Tall Rectangle Short Rectangle is equal to height Arm span is less than height. Arm span is greater than height For which graph, 1 or 2, is the line helpful in classifying a student's body shape as square, tall rectangle, or short rectangle? Explain. (4 points) Part C: Using the best model for prediction, calculate the predicted arm span for a senior with a height of 165 centimeters. (3 points)(10 points)
Answer #1

(a) there is positive association/correlation between height and arm span, points are concentrated towards III quadrant to I quadrant.

(b) Graph 1 is helpful, for this the regression equation is y^=64.494+0.5984x .

as it is the least square line from the given data and from this equation we can have all the 3 groups, but from the y^=x, only square group can be modelled

(c) for height x=165, the arm span= y^=64.494+0.5984x*165=163.23